Spring is compressed to a length of 20 cm by a 20 kg object on a 20 degrees inclined surface with coefficient of kinetic friction 0.30. What is the initial velocity of the object moving up the plane? The spring constant is 980 N/m


1
Expert's answer
2021-10-21T16:36:09-0400

Given:

"x=0.2\:\rm m"

"m=20\:\rm kg"

"k=980\:\rm N/m"


The law of conservation of energy gives


"mv^2\/2=kx^2\/2"

Hence the speed of the object


"v=\sqrt{k/m}x\\
=\sqrt{980/20}*0.2=1.4\:\rm m/s"